OpenStudy (anonymous):

Solve the equation algebraically and confirm graphically . x(2x-9)=18

OpenStudy (ahsome):

Expand the brackets on the LHSS \[x(2x-9)=18\] Expand this \[2x^2-9x=18\] Move the 18 to the LHS \[2x^2-9x-18=0\] You can either factorise or use the quadratic equation. Graph it with paper or GeoGebra to see if you got it right.

OpenStudy (ahsome):

Do you understand, @BrittanyO3343?